Please help me solve this: 3x-y=11 x+y=5 in elimination and substitution method.
Answer by stanbon(75887)

3x-y=11 x+y=5 in elimination and substitution method.
-------------------------
Elimination:
3x-y =11
x +y =5
-----
Add and solve for "x":
4x = 16
x = 4
----
Solve for "y" using x + y = 5
4 + y = 5
y = 1
---------
Substitution:
3x-y =11
x +y =5
-----
Solve the bottom equation for "x":
x = 5-y
Substitute into the top equation to solve for "y":
3(5-y) - y = 11
15 - 3y -y = 11
-4y = -4
y = 1
----
solve for "x" using x = 5-y
x = 5-1 = 4
====================
cheers,
Stan H.
========================